OpenStudy (anonymous):

Please show how to integrate x/((1+x^2)^2)?

OpenStudy (anonymous):

Try using integration by substitution, with u = x^2.

OpenStudy (anonymous):

put x^2=t then 2xdx=dt .. i.e. obtained by differentiating both sides of the equation wrt x. =>xdx=dt/2 hence, integral of xdx/((1+x^2)^2) will become integral of dt/2((1+t)^2)=-1/(2(1+t)) substituting back x^2=t, we get the answer = -1/(2(1+x^2))

OpenStudy (anonymous):

yeah..thank you

OpenStudy (fj):

hey dude you have to use this:#1(1+x^2)=u #2(2x)dx=du #3(xdx)=du/2 #4 integ(x/((1+x^2)^2)=integ(du/2u^2) then youre final answer:1/2(1+x^2)

OpenStudy (anonymous):